OpenStudy (anonymous):

You are given functions f and g such that f(n)=O(g(n)). Is f(n)∗log2(f(n)c)=O(g(n)∗log2(g(n))) ? (Here c is some constant >0. You can assume that f and g are always bigger than 1. a) true b)false c)depends on c d) depends on f and g

OpenStudy (anonymous):

take some example of f and g and chk if fn*logfn(n)<=gn*loggn

OpenStudy (foolaroundmath):

I think this violates the honor code of Coursera, so no replies :) Although I can give a hint. Use the definition of f(n) = O(g(n)) i.e. \[f(n) \le c*g(n) \forall n \ge n_{0}\] try fiddling around with this, multiply same functions or same sides of an inequality. This should be sufficient to get you to the answer.